OpenStudy (anonymous):

Estimate the circumference. Use 3.14 or 22/7 to approximate π. A. 220 ft B. 440 ft C. 879 ft D. 1758 ft

OpenStudy (anonymous):

PLZ HELP, fan and medal to best answer.

OpenStudy (midhun.madhu1987):

Where is the circle and Radius?

OpenStudy (triciaal):

circumference = 2*pi*r using pi = 22/7 insert the value of r, the radius in ft and calculate